OpenStudy (pixiedust1):

A carpenter is inserting a 3-foot brace between two beams as shown in the drawing (below). How far from the vertical beam will the brace reach? Round to the nearest tenth.

OpenStudy (pixiedust1):



OpenStudy (pixiedust1):

@mathstudent55

OpenStudy (pixiedust1):

@TheSmartOne

OpenStudy (pixiedust1):

I'm guessing this is the pythagorean theorum to solve, right?

TheSmartOne (thesmartone):

Pythagorean Theorem \(\Large\bf a^2 + b^2 = c^2\) where a and b are the two legs, and c is the hypotenuse so, solve for x x^2 + 1^2 = 3^3

TheSmartOne (thesmartone):

x^2 + 1^2 = 3^2 ***

TheSmartOne (thesmartone):

Yes, you are correct on that guess :)

OpenStudy (pixiedust1):

So then 1^2+3^2=3.16, which would be rounded to 3.2?

OpenStudy (pixiedust1):

Or actually, would it then be 2.8 ft because the missing angle is side b and not the hypotenuse?

TheSmartOne (thesmartone):

2.8 is correct :)

OpenStudy (pixiedust1):

Thank you!! Have a great day :D